# Structured methods for parameter inference and uncertainty quantification for mechanistic models in the life sciences

### v1

@@ Results
The model exhibits limit cycle dynamics for the parameter values chosen, completing around three cycles during the time period observed (Figure 2). Both the basic and the structured method were able to successfully recover the correct parameter values from the observed data. The univariate likelihood profiles are almost identical for both methods and are unimodal, implying that the parameters are identifiable (Figure 3). The 95% confidence intervals (range of values above the dotted horizontal line in Figure 3) contained the true parameter value for all target parameters for both methods.
@@
Figure 4: Results for the SEIR model from: (a) the basic method; (b) the structured method. Each panel shows the solution under the actual parameter values (solid curves); the solution under the maximum likelihood estimate for the parameter values (dashed curves); and the simulated data (dots). Where the dashed curve representing the solution at the MLE is not visible, this is because it coincides with the solid curve representing the solution at the true parameter values.
Figure 5: Normalised likelihood profiles for inferred parameters of the SEIR model: basic reproduction number ($R_{0}$); recovered to susceptible rate ($w$); observation probability ($p_{\mathrm{obs}}$); negative binomial dispersion parameter for observed data ($k$). The blue curves are from the basic method; the red curves are from the structured method; dashed vertical lines indicate the actual values (black) and the maximum likelihood estimates for each parameter under the basic method (blue) and structured method (red). Where the blue curve is not visible, this is because it coincides exactly with the red curve. Dotted horizontal line shows the threshold normalised log likelihood for the 95% confidence interval.
SEIR model
@@
When run on $M=500$ independently generated datasets, both methods again had similar accuracy, with a relative error in the MLE of 2.6% [1.3%, 4.9%] for the basic method, compared to 2.6% [1.2%, 4.6%] for the structured method (Table 1). The coverage rates were somewhat below 95%, and were slightly better for the structured method (90–93%) compared to the basic method (86–90%). However, the main advantage of the structured method was its efficiency, with a 44.5% [42.3%, 46.3%] reduction in the total number of calls required compared to the basic method. As for the predator-prey model, most of the efficiency improvement was in the profile likelihood component, although there was also a reduction of 9.2% [-2.6%, 19.5%] in the number of calls required for the MLE.
@@
When run on $M=500$, independently generated datasets, the relative error was 11.0% [6.6%, 17.3%] for both methods (Table 1). Coverage rates were reasonable (88–91% for the both methods). The structured method required 54.4% [51.2%, 56.7%] fewer calls to the model than the basic method did. Even for the MLE component alone, the structured method provided a substantial improvement in efficiency with 42.9% [33.9%, 50.3%] fewer calls than the basic method.
@@ Discussion
We have developed an adapted method for likelihood-based parameter inference and uncertainty quantification. The method is based on standard maximum likelihood estimation and profile likelihood, but exploits known structure in the mechanistic model to split a potentially high-dimensional optimisation problem into a nested pair of lower-dimensional problems. Only the outer problem requires the forward model to be solved at each step. Like standard maximum likelihood estimation, our method can be applied with a range of different noise models, provided a likelihood function exists.
@@
We have illustrated the new method on three toy models from different areas of the life sciences: a predator-prey model; a compartment-based epidemic model; and an advection-diffusion model. We have shown that the standard method and the new structured method provided comparable levels of accuracy in parameter estimates and coverage of inferred confidence intervals. However, the structured method was consistently more efficient, requiring substantially fewer calls to the forward model. In more complex models that are computationally expensive to solve, this would provide a major improvement in computation time.
@@
Although we have illustrated our method on three simple models, each with a single inner parameter, we anticipate there will be a broad class of models for which our approach could be applied. These include models where one or more dependent variables are linear in one or more parameters. This was the situation in the first two of our case studies. A similar situation would arise if observed model components were linear in an initial or boundary condition, or in a model with one-way coupling from a nonlinear to a linear component. Our approach is also applicable to models where a parameter effectively rescales one of the independent variables (typically space or time), which was the situation in the third case study. More generally, models which possess some sort of symmetry or invariance may also be candidates for applying our approach. These include travelling wave solutions (such as in models of a spreading population or other reaction–diffusion equations [29, 30]), similarity solutions (such as in models of chemotaxis [31, 32]), or scale invariance (such as in size-spectrum models of marine ecosystem dynamics [33, 34]).
@@
The process of specifying inner parameters and identifying the transformation relationship will in general be model-dependent. In models where such relationships exist, they may be revealed by a non-dimensionalisation procedure, which is a well-established technique in mathematical modelling [35]. A clear example of this would include a broad class of reaction–diffusion models with a logistic source term with carrying capacity $K$, where changing the value of $K$ linearly rescales the solution of the model [36].
@@
We have examined case studies in which data come from some specified noise process applied to the solution of an underlying deterministic model. Profile likelihood has also been applied to stochastic models, for example a stochastic model of diffusion in heterogeneous media [1].
In this first presentation of our structured inference approach, we have naturally focused on a range of applications where all parameters are identifiable. However, it will also be possible to apply our methodology to non-identifiable problems. Profile likelihood-based approaches are known to perform reasonably well in the face of non-identifiability whereas other approaches can give misleading results [8]. Another feature of our study is that we have applied our methods to standard cases where we treat the available data as fixed, but we note that we could adapt our method to deal with cases where inference and identifiability analysis is dynamically updated as new data becomes available [37].
@@
Finally, we remark that the work here has been presented in a frequentist, likelihood-based framework for parameter estimation and parameter identifiability analysis. We chose to work in a frequentist framework for the sake of algorithmic simplicity and computational efficiency. However, the results presented here could have been considered within a Bayesian setting by using the relevant likelihood functions within a sampling as opposed to an optimisation approach [38, 39]. Previous direct comparisons for parameter estimation and parameter identifiability suggest that taking a Bayesian approach leads to comparable results to a frequentist likelihood–based approach [8], albeit incurring an increased computational overhead since sampling is, in general, more computationally expensive than optimisation.
Our structured inference approach could, in principle, be used in a Bayesian setting by sampling from the distribution of outer parameters and, for each sample point, optimising the inner parameters using a likelihood or approximate likelihood function. Implementing this is beyond the scope of this paper but would be an interesting aim for future work.
@@ Acknowledgements
MJS is supported by the Australian Research Council (DP230100025). MJP acknowledges travel support from the Australian Research Council (DP200100177) The authors are grateful to the organisers of the 3rd New Zealand Workshop on Uncertainty Quantification and Inverse Problems, held at the University of Canterbury in 2023.

### v3

@@ Results
The model exhibited limit cycle dynamics for the parameter values chosen, completing around three cycles during the time period observed (Figure 2). Both the basic and the structured method were able to successfully recover the correct parameter values from the observed data. The univariate likelihood profiles were almost identical for both methods and were unimodal, implying that the parameters are identifiable (Figure 3). The 95% confidence intervals (range of values above the dotted horizontal line in Figure 3) contained the true parameter value for all target parameters for both methods. Obtaining the same results from both methods was an expected result, which provided confirmation that the structured method was working correctly.
@@
Figure 4: Results for the SEIRS model from: (a) the basic method; (b) the structured method. Each panel shows the solution under the actual parameter values (solid curves); the solution under the maximum likelihood estimate for the parameter values (dashed curves); and the simulated data (dots). Where the dashed curve representing the solution at the MLE is not visible, this is because it coincides with the solid curve representing the solution at the true parameter values.
Figure 5: Normalised likelihood profiles for inferred parameters of the SEIRS model: basic reproduction number ($R_{0}$); recovered to susceptible rate ($w$); observation probability ($p_{\mathrm{obs}}$); negative binomial dispersion parameter for observed data ($k$). The blue curves are from the basic method; the red curves are from the structured method; dashed vertical lines indicate the actual values (black) and the maximum likelihood estimates for each parameter under the basic method (blue) and structured method (red). Where the blue curve is not visible, this is because it coincides exactly with the red curve. Dotted horizontal line shows the threshold normalised log likelihood for the 95% confidence interval.
SEIRS model
@@
When run on $M=500$ independently generated datasets, both methods again had similar accuracy, with a relative error in the MLE of 2.6% [1.3%, 4.9%] for the basic method, compared to 2.6% [1.2%, 4.6%] for the structured method (Table 1). The coverage rates were slightly better for the structured method (90–93%) compared to the basic method (86–90%). However, the main advantage of the structured method was its efficiency, with a 44.5% [42.3%, 46.3%] reduction in the total number of calls required compared to the basic method. As for the predator-prey model, most of the efficiency improvement was in the profile likelihood component, although there was also a reduction of 9.2% [-2.6%, 19.5%] in the number of calls required for the MLE.
@@
When run on $M=500$ independently generated datasets, the relative error was 11.0% [6.6%, 17.3%] for both methods (Table 1). Coverage rates were reasonable (88–91% for both methods). The structured method required 54.4% [51.2%, 56.7%] fewer calls to the model than the basic method did. Even for the MLE component alone, the structured method provided a substantial improvement in efficiency with 42.9% [33.9%, 50.3%] fewer calls than the basic method.
@@ Discussion
We have developed an adapted method for likelihood-based parameter inference and uncertainty quantification. The method is based on standard maximum likelihood estimation and profile likelihood, but exploits known structure in the mechanistic model to split a high-dimensional optimisation problem into a nested pair of lower-dimensional problems using a similar approach to [17]. Only the outer problem requires the forward model to be solved at each step. Like standard maximum likelihood estimation, our method can be used with a range of different types of mathematical models and noise models, provided a likelihood function exists.
@@
We have illustrated the new method on three toy models from different areas of the life sciences: a predator-prey model; a compartment-based epidemic model; and an advection-diffusion PDE model. We have shown that the standard method and the new structured method provided comparable levels of accuracy in parameter estimates and coverage of inferred confidence intervals. However, the structured method was consistently more efficient, requiring substantially fewer calls to the forward model. In more complex models that are computationally expensive to solve, this would provide a major improvement in computation time.
@@
Although we have illustrated our method on three simple models, each with a single inner parameter, we anticipate there will be a broad class of models for which our approach could be applied. These include models where one or more dependent variables are linear in one or more parameters, as considered by [17, 18]. This was the situation in the first two of our case studies. A similar situation would arise if observed model components were linear in an initial or boundary condition, or in a model with one-way coupling from a nonlinear to a linear component. Our approach is also applicable to models where a parameter effectively rescales one of the independent variables (typically space or time), which was the situation in the third case study. More generally, models which possess some sort of symmetry or invariance may also be candidates for applying our approach. These include travelling wave solutions (such as in models of a spreading population or other reaction–diffusion equations [33, 34]), similarity solutions (such as in models of chemotaxis [35, 36]), or scale invariance (such as in size-spectrum models of marine ecosystem dynamics [37, 38]). Exploring the range of models and classes of parameter relationships for which our methodology is applicable is an avenue for further research.
@@
The process of specifying inner parameters and identifying the transformation relationship will in general be model-dependent. In models where such relationships exist, they may be revealed by a non-dimensionalisation procedure, which is a well-established technique in mathematical modelling [39]. A clear example of this would include a broad class of reaction–diffusion models with a logistic source term with carrying capacity $K$, where changing the value of $K$ linearly rescales the solution of the model [40]. However, we acknowledge that not all models will be amenable to the structured inference we have presented because the relationship between parameters of interest and model solution will be unknown or too complex to identify a priori.
@@
We have examined case studies in which data come from some specified noise process applied to the solution of an underlying deterministic model. In the second and third case studies, where the noise distribution had a variance parameter, we did not assume this was known a priori, but included it as a target for inference for both the basic and the structured method. Profile likelihood has also been applied to stochastic models, for example a stochastic model of diffusion in heterogeneous media [1]. Another feature of our study is that we have applied our methods to standard cases where we treat the available data as fixed, but we note that we could adapt our method to deal with cases where inference and identifiability analysis is dynamically updated as new data becomes available [41]. How the size of the efficiency improvement scales with the number of inner and outer parameters in more complex models is also an interesting question for future work.
@@
All results in this study focus on assessing practical identifiability since we are interested in real-world problems, which involve working with imperfect, noisy and sparse data. Alternatively, we could also consider the structural identifiability, which considers the case where we have highly-idealised, noise-free observations [42]. Structural identifiability for ODE models can be assessed using algebraic methods that are available in several software packages, such as GenSSI [43], DAISY [44] and STRIKE-GOLDD [45]. Assessing practical identifiability is a stronger condition than assessing structural identifiability because structurally identifiable models can turn out to be practically non-identifiable when working with finite, noisy data. Previous work has established that likelihood-based methods can perform well regardless of whether a model is structurally or practically non-identifiable [46, 47].
We have worked with a likelihood-based framework, which we chose for the sake of algorithmic simplicity and computational efficiency. In particular, working with profile likelihood is often faster than sampling based methods, such as MCMC [8], and this is particularly relevant for poorly identified problems. All problems we have considered here involve parameters that are well-identified by the data. We note that if the variance in the noise was larger than assumed here, or if the data were sparser, this could result in parameters being poorly identified. We anticipate that, where identifiability issues exist, they would affect both the basic and the structured methods similarly. It is possible that the structured approach might speed up the computations needed to diagnose the identifiability issue. However, we leave a detailed exploration of this for future work. One way of doing this would be to seek an appropriate re-parameterisation of the log-likelihood function, which we leave for future consideration.
Our structured inference approach could, in principle, be used in a Bayesian setting by sampling from the distribution of outer parameters and, for each sample point, optimising the inner parameters using a likelihood or approximate likelihood function (see also [48]). Implementing this is beyond the scope of this paper but would be an interesting aim for future work.
@@ Acknowledgements
MJS is supported by the Australian Research Council (DP230100025). MJP acknowledges travel support from the Australian Research Council (DP200100177). The authors thank the organisers of the 3rd New Zealand Workshop on Uncertainty Quantification and Inverse Problems, held at the University of Canterbury in 2023. The authors are grateful to five anonymous reviewers for comments on previous versions of this manuscript.